OpenStudy (g0thw1ng):

When four times a number is added to 8 times the number, the result is 36. What is the number?

OpenStudy (fuzzpuppy2):

ok so i would start out by taking the 8 times a number and the four times a number and putting them together to get 12 times a number

OpenStudy (fuzzpuppy2):

8 x a number = 36 so the opposite of times is division

OpenStudy (fuzzpuppy2):

oops i meant 12 x a number = 36

OpenStudy (g0thw1ng):

I keep getting 4.5 and-4.5

OpenStudy (fuzzpuppy2):

ok so re write it so that the question now is 36 / 12 = a number what do you get?

OpenStudy (g0thw1ng):

3

OpenStudy (fuzzpuppy2):

so to check your answer plug it in. 4 x 3 = 12 + 8 x 3 = 24

OpenStudy (fuzzpuppy2):

12 + 24 = 36 so your answer would be 3

OpenStudy (fuzzpuppy2):

i hope this helped you! im still in middle school though and since this is a "college" question im not sure if its 100% right. But taking the information you gave me im pretty sure i did it right. again i hope this helped, and if not im sorry -.-

OpenStudy (g0thw1ng):

Don't worry you guided me to the right direction where I was lost thanks